OpenStudy (anonymous):

solve (15(0.95)^(n+1))=20(0.94)^(n),n

OpenStudy (mathstudent55):

Is this the equation? \(\large 15(0.95)^{n + 1} = 20 (0.94^n) \)

OpenStudy (anonymous):

Yes

OpenStudy (mathstudent55):

First, divide both sides by 20.

OpenStudy (anonymous):

It okay, I solved it

OpenStudy (mathstudent55):

\(\large 0.75(0.95)^{n + 1} = 0.94^n\) \(\large 0.75(0.95^n)(0.95) = 0.94^n\) \(\large (0.95)(0.75)(0.95^n) = 0.94^n\) \( \large( \dfrac{0.95}{0.94} \large)^n = 0.95^{-1}0.75^{-1} \) \(\large \ln (\dfrac{0.95}{0.94}\large)^n = \ln (0.95^{-1}0.75^{-1} \)) \(n\ln \dfrac{0.95}{0.94} = -\ln 0.95 - \ln 0.75 \) \(n( \ln 0.95 - \ln 0.94) = -\ln 0.95 - \ln 0.75 \) \(n = \dfrac{-\ln 0.95 - \ln 0.75}{ \ln 0.95 - \ln 0.94 } \)

OpenStudy (mathstudent55):

\(n = -\dfrac{\ln 0.95 + \ln 0.75}{ \ln 0.95 - \ln 0.94 }\) \(n \approx 32.033\)